OpenStudy (josee):

Limit x->7+ to (x-8)/(x-7)?

jimthompson5910 (jim_thompson5910):

First, let's set up a table. To create the table, simply plug in values of x and record the corresponding values of y. For instance, when x = 6, then y = (6-8)/(6-7) = -2/(-1) = 2, so one line is 6 2 So do this for values near 7 to get x y 6.00000 2.00000 6.10000 2.11111 6.20000 2.25000 6.30000 2.42857 6.40000 2.66667 6.50000 3.00000 6.60000 3.50000 6.70000 4.33333 6.80000 6.00000 6.90000 11.00000 7.00000 undefined 7.10000 -9.00000 7.20000 -4.00000 7.30000 -2.33333 7.40000 -1.50000 7.50000 -1.00000 7.60000 -0.66667 7.70000 -0.42857 7.80000 -0.25000 7.90000 -0.11111 8.00000 0.00000 Notice how as x decreases towards x=7, the y values are becoming larger and larger in the negative direction. If you keep going, these values will approach negative infinity. So \[\large \lim_{x\to7^{+}}\left(\frac{x-8}{x-7}\right)=-\infty\]

OpenStudy (josee):

Ohhhhhhhh! What if 7 is coming from the left? As x approaches 7-

jimthompson5910 (jim_thompson5910):

then we look at the table for x values that are slowly increasing to x=7 these values point to y values that are increasing so \[\LARGE \lim_{x\to7^{-}}\left(\frac{x-8}{x-7}\right)=\infty\]

OpenStudy (josee):

Omg you taught me so much right now.. and since when x comes from the left is infinity and as it comes from the right its negative infinity there isnt a limit as x just apporaches 7... :D

jimthompson5910 (jim_thompson5910):

exactly, the limit \[\large \lim_{x\to7^{-}}\left(\frac{x-8}{x-7}\right)\] doesn't exist because \[\LARGE \lim_{x\to7^{-}}\left(\frac{x-8}{x-7}\right)\neq \lim_{x\to7^{+}}\left(\frac{x-8}{x-7}\right)\]

OpenStudy (josee):

Thank you so much:)

OpenStudy (josee):

\[\lim_{x \rightarrow 0}(3/x^2+3x)-(1/x)\]?